OpenStudy (anonymous):

Juan jogs each day from his house to a local park, and then he jogs back along the same route. On his way to the park, he averages 4 miles per hour. On his way home, he averages 6 miles per hour. If the total trip takes 1 1/2 hours, which equation can be used to find x, the distance from Juan’s house to the park? x/4+x/6=1 x/4+x/6=1 1/2 4x6x=1 1/2 4x+6x=1

OpenStudy (anonymous):

speed 4 miles /hr means he covers 4 miles in an hour covers 1 mile in =1/4 hr covers x miles in=(1/4 )*x hrs=x/4 hrs for return journey speed 6 miles/hr means he covers 6 miles in one hour covers 1 mile in=1/6 hrs covers x miles in=(1/6)*x hrs Total time taken=x/4 +x/6=?